# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 3 Compiled Bell inequalities
The marginal distribution $p(a|x)$ is independent of the second input $y$ due to the sequential nature of the protocol. Later, we will need to establish a type of independence between the second outcome $b$ and the first input $x$. To do so we will need to consider the distributions of the decrypted outputs and appeal to the security promise of the QHE scheme. Specifically, we require a key lemma which has appeared in several works [NZ23, CMM+24]. We borrow a version from [BVB+24], which we refer the reader to for proof.
Lemma 3.1 ([BVB+24], Lemma 7).
@@
Let $\lambda\in\mathbb{N}$ be the security parameter for a QHE scheme. There exists a negligible function $\delta_{\mathsf{QHE}}$ of $\lambda$ such that the following holds. For any two QPT sampleable distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ over plaintext inputs $\mathcal{A}$, QPT preparable state $|\Psi\rangle$, and QPT implementable operation $\mathcal{N}$, it holds that
@@
$\left|\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{1}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha,\beta}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle-\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha,\beta}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle\right|\leq\delta_{\mathsf{QHE}}.$
@@
The approximate no-signalling conditions from Alice to Bob can then be seen by applying Lemma 3.1, since
@@
$\Big{|}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle-\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x^{\prime}}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle\Big{|}\leq\delta_{\mathsf{QHE}}$ (3.5)
@@ 3.1 Quantum bounds for compiled inequalities
A compiled (quantum) model describes the correlation $\{p(a,b,|x,y)\}$ observed in a compiled Bell scenario. A compiled Bell functional is a linear functional $I$ evaluated on correlations realized by compiled models. That is
@@
$I=\sum_{abxy}w_{abxy}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ \chi:\mathsf{Dec}(\chi)=x\end{subarray}}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle.$ (3.6)
The compilation procedure preserves the classical bound of the Bell inequality (up to negligible error)[KLVY23]. Hence in the compiled setting, classical provers cannot effectively violate a Bell inequality by more than they could in the classical (bipartite) scenario.
@@
We now turn our attention to quantum provers in the compiled setting. The results of [KLVY23] imply that a quantum prover in the compiled setting can achieve the maximal bound in the bipartite setting. However, the existence of a quantum compiled behaviour which exceeds the maximal quantum bell violation in the bipartite case (by more than negligible factors) has not been ruled out. Nonetheless, in several cases (like the CHSH inequality and more generally all XOR games [CMM+24]) we know that the quantum compiled behaviour cannot exceed the value $\eta^{\mathrm{Q}}$ by more than negligible amounts. One technique for establishing such bounds was introduced in [NZ23] and uses SOS techniques to bound the quantum violation of the compiled Bell functional.
3.2 Extended definition of the pseudo-expectation
@@
Our approach builds off the methods used in [NZ23] and [CMM+24] using pseudo-expectations. This SOS approach necessitates the definition of a pseudo-expectation; defined as a unital, linear map from a subspace $\mathcal{T}$ of the algebra generated by $\{M_{a|x},N_{b|y}\}_{a\in\mathcal{A},x\in\mathcal{X},b\in\mathcal{B},y\in\mathcal{Y}}$ to the complex numbers $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}:\mathcal{T}\mapsto\mathbb{C}$. In the case $n=m=2$, it suffices to define the pseudo expectation $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the observables $A_{x}=\sum_{a\in\{0,1\}}(-1)^{a}M_{a|x},\ B_{y}=\sum_{b\in\{0,1\}}(-1)^{b}N_{b|y}$ and require that they are mapped to their expectations in the compiled scenario [^4] Though in the following we define $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ for $n=m=2$, this can be directly extended to arbitrary Bell scenarios by defining $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the POVM elements $M_{a|x},N_{b|y}$ in an analogous way.. We also require $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ to be approximately positive on squares; that is, for all $w\in\mathcal{T}$, $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[w^{\dagger}w]\geq-\delta$ for some $\delta\geq 0$. In previous works, the definition of the pseudo-expectation has been restricted to at most degree 2 monomials, consisting of at most one Alice and one Bob observable:
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}B_{y}]$ $:=\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$ (3.7)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}A_{x^{\prime}}]$ $:=\delta_{x,x^{\prime}},$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}B_{y^{\prime}}]$ $:=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}B_{y^{\prime}}|\Psi_{\alpha|\chi}\rangle,$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}]$ $:=\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|\Psi_{\alpha|\chi}\rangle$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}]$ $:=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\mathbbm{1}]$ $:=1,$
@@
where $\mathop{\mathbb{E}}_{x\in\mathcal{X}}$ denotes the expectation according to an arbitrary fixed distribution over $\mathcal{X}$. This is already sufficient to handle known SOS decompositions for a variety of well-studied Bell inequalities whenever the polynomials are expressed in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x,y}$. However, there are Bell inequalities, such as the tilted-CHSH inequality, for which no known SOS decomposition exists in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x,y}$.
@@
The contribution of this section is to expand the definition of the pseudo-expectation to the basis encompassing all monomials in $A_{x},B_{0},B_{1}$, for a fixed $x\in\mathcal{X}$. This allows us to handle more general SOS decompositions, and in particular, the tilted-CHSH inequalities. Let $w(A_{x},B_{0},B_{1})$ be a monomial in the elements $\{A_{x},B_{0},B_{1}\}$. Importantly, $x$ is fixed, and we do not consider monomials of the form $A_{0}A_{1}B_{y}$. Let $\bar{w}$ be the canonical form of $w$ under the relations $[A_{x},B_{y}]=0$, $(B_{y})^{2}=(A_{x})^{2}=\mathbbm{1}$, where all $A_{x}$ terms are commuted to the left. Since we only consider one value of $x$, these will all be of the form $(A_{x})^{i}\bar{w}(B_{0},B_{1})$ for some $i\in\{0,1\}$, where the monomial $\bar{w}(B_{0},B_{1})$ cannot be reduced further. We then define the pseudo-expectation
@@
From the above definitions, we state the main result of this section in the following lemma, which can be applied generally to any polynomial expressible in the basis $\{A_{x},B_{0},B_{1}\}$.
Lemma 3.2.
Let $A_{x}$ and $\{B_{y}\}_{y\in\mathcal{Y}}$ be binary observables, and let

### v3

@@ 3 Compiled Bell scenarios
Note that the marginal distribution $p^{(\lambda)}(a|x)$ obtained from Eq. 3.4 will be independent of the second input $y$ due to the sequential nature of the protocol. However, the marginal $p^{(\lambda)}(b|y,x)$ currently depends on $x$. The aim of what follows is to establish a computational independence between this distribution and the inputs $x$. To do so we will need to consider the distributions of the decrypted outputs and appeal to the security promise of the QHE scheme. Specifically, we require a key lemma which has appeared in several works [NZ23, CMM+24, BVB+24]. We borrow a version from [KMP+24] and we refer the reader to the reference for the proof.
Lemma 3.1 ([KMP+24], Proposition 4.6).
@@
Let $\widetilde{\mathcal{\mathrm{Q}}}^{(\lambda)}$ be a compiled quantum model, and $\mathcal{N}^{(\lambda)}=w(\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B},y\in\mathcal{Y}})$ be a monomial in the measurement operators $\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B},y\in\mathcal{Y}}$, where $\lambda\in\mathbb{N}$ is the security parameter for a fixed QHE scheme. Then, for any two QPT sampleable distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ over plaintext inputs $x\in\mathcal{X}$ there exists a negligible function $\mathrm{negl}(\lambda)$ of the security parameter $\lambda$ such that the following holds
@@
$\left|\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{1}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\chi:\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}^{(\lambda)}|\mathcal{N}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle-\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\chi:\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}^{(\lambda)}|\mathcal{N}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\right|$
$\leq\mathrm{negl}(\lambda).$
@@
The approximate no-signalling conditions from Alice to Bob can then be seen by applying Lemma 3.1 to the monomials of degree 1 in the QPT measurement operators $\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B},y\in\mathcal{Y}}$, since
@@
$\Big|\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}^{(\lambda)}|N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle-\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x^{\prime})=\chi}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}^{(\lambda)}|N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\Big|\leq\mathrm{negl}(\lambda)$ (3.5)
@@
In the above statements, the measurements are completely general, and the states are sub-normalized vectors. The following lemma shows that when considering the compiled value, we can assume that the states and measurement operators in the compiled strategy are pure and projective.
Lemma 3.2.
Let $\mathcal{H}^{{}^{\prime}(\lambda)}$ be the Hilbert space of the device, and $\{\{\rho_{\alpha|\chi}^{(\lambda)}\}_{\alpha\in\bar{\mathcal{A}}}\}_{\chi\in\bar{\mathcal{X}}}$ be a family of QPT-preparable sub-normalized states on $\mathcal{H}^{{}^{\prime}(\lambda)}$ after the first round. Let $\{\{N_{b|y}^{{}^{\prime}(\lambda)}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}$ be a family of QPT-implementable POVMs on $\mathcal{H}^{{}^{\prime}(\lambda)}$, which induce the behaviour $p^{(\lambda)}(\alpha,b|\chi,y)={\mathrm{t}r}[N_{b|y}^{{}^{\prime}(\lambda)}\rho_{\alpha|\chi}^{(\lambda)}]$. Then there exists a Hilbert space $\mathcal{H}^{(\lambda)}$, a family of QPT-preparable sub-normalized states $\{\{|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\}_{\alpha\in\bar{\mathcal{A}}}\}_{\chi\in\bar{\mathcal{X}}}$ in $\mathcal{H}^{(\lambda)}$, and a family of QPT-implementable PVMs $\{\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}$ on $\mathcal{H}^{(\lambda)}$ which satisfy
$\langle\Psi_{\alpha|\chi}^{(\lambda)}|N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle=p^{(\lambda)}(\alpha,b|\chi,y),\ \ \forall\alpha\in\bar{\mathcal{A}},\,\chi\in\bar{\mathcal{X}},\,b\in\mathcal{B},\,y\in\mathcal{Y}.$ (3.6)
See Section A.2 for the proof of Lemma 3.2.
We say a compiled model $\widetilde{\mathcal{\mathrm{Q}}}^{(\lambda)}=(\widetilde{\mathcal{H}}^{(\lambda)},\{|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}},\{\{N_{b|y}^{(\lambda)}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}})$ is pure and projective whenever the states $|\Psi_{\alpha|\chi}^{(\lambda)}\rangle$ are all pure and the measurements $N_{b|y}^{(\lambda)}$ are all projective (i.e. PVMS).
@@ 3.1 Quantum bounds for compiled inequalities
A compiled (quantum) model $\widetilde{\mathcal{\mathrm{Q}}}^{(\lambda)}$ describes the correlations $\mathsf{p}^{(\lambda)}=\{p^{(\lambda)}(a,b|x,y)\}_{a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}}$ observed in a compiled Bell scenario. A compiled Bell functional is a linear functional $I^{(\lambda)}$ evaluated on correlations realized by compiled models. That is
@@
$I^{(\lambda)}=\sum_{abxy}w_{abxy}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ \chi:\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi^{(\lambda)}_{\alpha|\chi}|N^{(\lambda)}_{b|y}|\Psi^{(\lambda)}_{\alpha|\chi}\rangle$. (3.7)
By the properties of the compilation procedure [KLVY23, Theorem 3.2], Bell inequalities are preserved under compilation (up to negligible error). In particular, for large security parameter, efficient classical provers cannot violate a Bell inequality by much more than they could in the (bipartite) scenario. From now on, we will suppress the security parameter $\lambda\in\mathbb{N}$ along with the expectation over secret keys $\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}$ and simply write the expectation for a fixed key. In particular, we express the compiled model as $\widetilde{\mathcal{\mathrm{Q}}}$ and Eq. 3.7 as
$I=\sum_{abxy}w_{abxy}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle.$
@@
We now turn our attention to the maximum value $I$ can take in the compiled setting with an efficient quantum prover. The results of [KLVY23] imply that an efficient quantum prover can achieve the same violation in the bipartite setting. However, the existence of a quantum compiled behavior which exceeds the maximal quantum Bell violation in the bipartite case (by more than negligible factors) has not been ruled out. Nonetheless, in several cases (like the CHSH inequality and more generally all XOR games [CMM+24]) we know that the quantum compiled behavior cannot exceed the value $\eta^{\mathrm{Q}}$ by more than negligible amounts. One technique for establishing such bounds was introduced in [NZ23] and uses SOS techniques to bound the quantum violation of the compiled Bell functional.
3.2 Extending the pseudo-expectations
@@
Our approach builds off the methods used in [NZ23] and [CMM+24]. To explain this approach we recall that a pseudo-expectation is a unital, linear map from a subspace $\mathcal{T}$ of the algebra generated by $\{M_{a|x},N_{b|y}\}_{a\in\mathcal{A},x\in\mathcal{X},b\in\mathcal{B},y\in\mathcal{Y}}$ to the complex numbers, $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}:\mathcal{T}\to\mathbb{C}$, which is determined by a compiled quantum model $\widetilde{\mathcal{\mathrm{Q}}}$. In the case $n=m=2$, it suffices to define the pseudo-expectation $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the observables $A_{x}=\sum_{a\in\{0,1\}}(-1)^{a}M_{a|x},\ B_{y}=\sum_{b\in\{0,1\}}(-1)^{b}N_{b|y}$ and require that they are mapped to their expectations in the compiled scenario [^2] Though in the following we define $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ for $n=m=2$, this can be directly extended to arbitrary Bell scenarios by defining $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the POVM elements $M_{a|x},N_{b|y}$ in an analogous way.. We further assume that all measurements are projective (cf. Eq. 3.6). In previous works, the definition of the pseudo-expectation had been restricted to monomials consisting of at most one Alice and one Bob observable as outlined below:
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}B_{y}]$ $=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$ (3.8)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}A_{x^{\prime}}]$ $=\delta_{x,x^{\prime}},$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}B_{y^{\prime}}]$ $=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}B_{y^{\prime}}|\Psi_{\alpha|\chi}\rangle,$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}]$ $=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|\Psi_{\alpha|\chi}\rangle,$
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}]$ $=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\mathbbm{1}]$ $=1,$
@@
where $\mathop{\mathbb{E}}_{x\in\mathcal{X}}$ denotes the expectation according to an arbitrary fixed distribution over $\mathcal{X}$. This is already sufficient to handle known SOS decompositions for a variety of well-studied Bell inequalities whenever the polynomials are expressed in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x\in\mathcal{X},y\in\mathcal{Y}}$. However, there are Bell inequalities, such as the tilted-CHSH inequality [BP15, BSB24], for which no known SOS decomposition exists in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x\in\mathcal{X},y\in\mathcal{Y}}$.
@@
The contribution of this section is to expand the definition of the pseudo-expectation to the basis encompassing all monomials in $A_{x},B_{0},B_{1}$, for a fixed $x\in\mathcal{X}$, in a way that is approximately non-negative on Hermitian squares. This allows us to handle more general SOS decompositions, and in particular, the tilted-CHSH inequalities. Let $w(A_{x},B_{0},B_{1})$ be a monomial in the elements $\{A_{x},B_{0},B_{1}\}$. Importantly, $x$ is fixed, and we do not consider monomials of the form $A_{0}A_{1}B_{y}$ for example. Let $\bar{w}$ be the canonical form of $w$ under the relations $A_{x},B_{y}]=0$, $(B_{y})^{2}=(A_{x})^{2}=\mathbbm{1}$, where all $A_{x}$ terms are commuted to the left. Since we only consider one value of $x$, these will all be of the form $(A_{x})^{i}\bar{w}(B_{0},B_{1})$ for some $i\in\{0,1\}$, where the monomial $\bar{w}(B_{0},B_{1})$ cannot be reduced further. We then define the pseudo-expectation
@@
From the above definitions, we next state the main result of this section, which can be applied generally to any polynomial expressible in the basis $\{A_{x},B_{0},B_{1}\}$.
Theorem 3.3.
Let $\{A_{x}\}_{x\in\mathcal{X}}$ and $\{B_{y}\}_{y\in\mathcal{Y}}$ be binary observables, and let